# Topology-Driven Quantum Architecture Search Framework

### v1

@@ 4.3.1 TD-QuantumSupernet framework
In this section, we use QuantumSupernet to implement the TD-QAS framework, followed by a series of numerical simulations. The core algorithm of QuantumSupernet is described in Section 2.1.1. Therefore, we will first explain how to implement the TD and GT phases using QuantumSupernet. Next, we define the hyperparameters for various numerical simulation tasks, highlighting their impact on the comparison variables. Finally, we execute the algorithm using these hyperparameters and analyze the results of the numerical simulations. For the simulations, we select the same VQE and classification tasks used in QuantumSupernet.
Overview of Framework
@@
In this section, we provide an overview of the implementation process of the TD-QAS framework using QuantumSupernet. The TD phase begin by defining the topology search space, which is determined by the structure of the supernet and the native gate set. The number of qubits ($N_{\text{qubit}}$) and the number of layers ($N_{\text{layer}}$) are specified by the task. Additionally, the native gate set is set to ${R_{x},I,CNOT,CI}$, where $CI$ refers to the Controlled-Identity Gate. If a sampled topology instance contains an $I$ or $CI$ gate at a specific location, it indicates that no quantum gate will be placed at that position, and the GT phase will skip fine-tuning for that location. Next, a randomly searching algorithm is exploring potential high-performance topologies. The performance of each topology instance is evaluated using shared parameters, with feedback used to update these parameters. Specifically, we use $N_{\text{experts}}$ shared parameters to assess the performance of each topology, which allows for more accurate performance evaluations. The shared parameters are trained over $N_{\text{train}}$ iterations. During the first $T_{\text{warm}}$ iterations, a randomly selected expert from the $N_{\text{experts}}$ pool evaluates the topology instances and updates the optimal shared parameters. After the warm-up phase, all $N_{\text{experts}}$ participate in the evaluation, but only the shared parameters with the best performance are updated. After the convergence of shared parameters, we evaluate $N_{\text{search}}$ topology instances and select the most optimal topology. In the GT phase, the gate-type search space consists of the previously identified topology and the native gate set $A_{\text{g}}$. The shared parameters for the GT phase are inherited from the TD phase. This inheritance mechanism involves replicating the shared parameters from the TD phase based on the number of single-qubit and two-qubit gates available in the native gate pool. To enhance the accuracy of the evaluation, we perform additional training on the inherited parameters for a few extra iterations, denoted as $T_{\text{extra}}$, which further refines the assessment of specific circuit performances.
@@
With the designed hyperparameters, we calculate the comparison variables, including the search space size and the total quantum computational cost (QCC). Since the supernet consistently applies both single-qubit and two-qubit gates to every qubit, the original search space is calculated as $(|A[\text{single}]|^{N_{\text{qubit}}}\cdot|A[\text{double}]|^{N_{\text{qubit}}})^{N_{\text{layer}}}$, where $|A[\text{single}]|$ denotes the number of single-qubit gates in the native gate set. The topology and gate-type search space is then defined as $(2^{\text{qubit}})^{N_{\text{layer}}}$ and $|A[\text{single}]|^{N_{\text{layer}}-\text{x}}\cdot|A[double]|^{\text{x}}$, where $2^{N_{\text{qubit}}}$ represents the combinations of single and double placeholder gates across different qubits, and $x$ is the number of double placeholder gates in the searched topology. The quantum computational cost (QCC) for QuantumSupernet and the TD phase is given by $\sum_{\text{1}}^{T_{\text{warm}}}T_{\text{i}}+\sum_{T_{\text{warm}}}^{T}T_{\text{i}}*N_{\text{experts}}+\sum_{\text{1}}^{N_{\text{search}}}T_{\text{i}}*N_{\text{experts}}$ The QCC for the GT phase is given by $\sum_{\text{1}}^{N_{\text{search}}}T_{\text{i}}\cdot(1+T_{\text{extra}})$. Each circuit of quantum computational costs $T_{\text{i}}$ is defined as (7).
@@
In this section, we evaluate the performance of the TD-QAS framework through numerical simulations. We first replicate the ground state energy estimation task for a 4-qubit hydrogen molecule, which is consistent with the simulation in Ref. [23]. While this small-scale task provides valuable insight, it does not fully showcase the potential of the TD-QAS framework. Therefore, we extend the validation to larger tasks, including the 5-qubit Heisenberg model.
For the simulation task involving the estimation of the ground state energy of hydrogen (which has a ground state energy of -1.13618 Hartree), the construction of the Hamiltonian is detailed in QuantumSupernet [23]. Before performing the numerical simulations, we define the hyperparameters and compare the search space sizes of the QuantumSupernet and TD-QAS methods. Specifically, the number of qubits ($N_{\text{qubit}}$) is set by to 4 and the number of layers ($N_{\text{layer}}$) is to 3. For the QuantumSupernet, the native gate set $A$ is defined as $\{I,R_{y},R_{z},CNOT,CI\}$. In the TD-QAS framework, the gate sets are defined as $A_{\text{t}}=\{R_{y},I,CNOT,CI\}$ for the topology-driven phase and $A_{\text{g}}=\{R_{y},R_{z},CNOT\}$ for the gate tuning phase. Under these settings, the search space for the QuantumSupernet is calculated as $(3^{4}\cdot 2^{4})^{3}$. The search space for the TD phase is calculated as $(2^{4}\cdot 2^{4})^{3}$, while the GT phase operates within a search space ranging from 1 to $2^{12}$. This decoupling results in a search space reduction ratio of $2.5\times 10^{1}$. Due to the reduced search space, we use smaller values for the number of experts ($N_{\text{experts}}$) and the total number of iterations ($T$) to enhance efficiency. Specifically, for the QuantumSupernet, we set $N_{\text{experts}}=5$, $T=500$, $T_{\text{warm}}=200$, and $N_{\text{search}}=500$. In the TD phase, these parameters are adjusted to $N_{\text{experts}}=1$, $T=200$, $T_{\text{warm}}=100$, and $N_{\text{search}}=500$. In the GT phase, we set $T_{\text{extra}}=1$. Given the small problem size and native gate set, an exhaustive search is feasible, and thus $N_{\text{search}}$ is omitted in the GT phase.
@@
To further demonstrate the capabilities of the TD-QAS framework for larger-scale problems, we applied it to the ground state energy estimation of the 5-qubit Heisenberg model. The Hamiltonian for the Heisenberg model is defined as $\sum_{\text{i=1}}^{\text{n}}\left(X_{\text{i}}X_{\text{i+1}}+Y_{\text{i}}Y_{\text{i+1}}+Z_{\text{i}}Z_{\text{i+1}}\right)+\sum_{\text{i=1}}^{\text{n}}Z_{\text{i}}$ , and the ground state energy is -8.47213. The hyperparameters for this simulation are defined as follows: $N_{\text{qubit}}$=5, $N_{\text{layer}}$= 6, $A$={$I$, $R_{x}$, $R_{y}$, $R_{z}$, $CNOT$, $CI$}, $A_{\text{t}}$={$R_{y}$, $I$, $CNOT$, $CI$}, and $A_{\text{g}}$={$R_{x}$, $R_{y}$, $R_{z}$, $CNOT$}. With these settings, the search space for the QuantumSupernet is calculated as $(4^{5}\cdot 2^{5})^{6}=2^{80}$, while the search space for the TD and GT phases is $2^{60}$ and $1\sim 3^{30}$, respectively. This results in a search space reduction ratio of approximately $1\times 10^{7}$. To evaluate the performance of the TD-QAS framework and compare it to the QuantumSupernet, we executed simulations with different values of $N_{\text{experts}}$. Specifically, we set $N_{\text{experts}}=1$ and $N_{\text{experts}}=5$. The hyperparameters for the QuantumSupernet are $T=500$, $T_{\text{warm}}=200$, and $N_{\text{search}}=500$. For the TD-QAS framework, the parameters are adjusted as follows: $T=200$, $T_{\text{warm}}=100$, and $N_{\text{search}}=300$. In the GT phase, we set $T_{\text{extra}}=1$ and $N_{\text{search}}=100$.
@@
The results of the numerical simulations are presented in Table 2. With the same $N_{\text{experts}}$, the TD and GT methods outperform the QuantumSupernet. This suggests that the TD-QAS framework enhances the performance of the QuantumSupernet by efficiently reducing circuit complexity. When comparing simulations with different numbers of experts ($N_{\text{experts}}=1$ and $N_{\text{experts}}=5$), we observe that the QuantumSupernet shows a higher dependency on $N_{\text{experts}}$. This is likely due to the larger number of parameterized quantum gates in the circuits searched by QuantumSupernet, which results in greater sensitivity to the number of experts. In contrast, the circuits found by the TD and GT methods have fewer parameters, leading to more stable performance across varying values of $N_{\text{experts}}$. This highlights the TD-QAS framework’s ability to reduce the training burden and, consequently, the QCC. Additionally, the circuits discovered in both the TD and GT phases exist within the operation search space, but were not identified by the QuantumSupernet. This suggests that the TD-QAS framework may help mitigate the risk of the QAS algorithm getting trapped in local optima, thereby facilitating the discovery of more optimal quantum circuits.
@@
To assess the performance improvements of the TD-QAS framework on more complex tasks, we conducted a numerical evaluation using an 8-qubit quantum state classification task. Specifically, we employed a QNN to address a binary classification problem, where the input consisted of quantum states with varying levels of entanglement, characterized by Concurrence Entropy (CE = 0.15 and CE = 0.45). The quantum dataset used in this experiment is based on the dataset from Ref. [46].
The hyperparameter settings for the numerical simulation are detailed in follow, $N_{\text{qubit}}$=8, $N_{\text{layer}}$= 6, $A$={$I$, $R_{x}$, $R_{y}$, $R_{z}$, $CNOT$, $CI$}, $A_{\text{t}}$={$R_{y}$, $I$, $CNOT$, $CI$}, and $A_{\text{g}}$={$R_{x}$, $R_{y}$, $R_{z}$, $CNOT$}. With these parameters, we calculate that the search space for the QuantumSupernet is $2^{144}$, while the search space for the TD and GT phases is $2^{96}$ and $1\sim 3^{30}$, respectively. This results in a search space reduction ratio of approximately $2.6\times 10^{6}$. The simulation hyperparameters for the QuantumSupernet are set as $T=500$, $T_{\text{warm}}=200$, and $N_{\text{search}}=500$. For the TD-QAS framework, we set $T=200$, $T_{\text{warm}}=100$, and $N_{\text{search}}=300$ in the TD phase, while in the GT phase, $T_{\text{extra}}=1$ and $N_{\text{search}}=100$.
Table 3 summarizes the simulation results for the state classification task. The TD-QAS framework successfully identifies higher-performance circuits compared to the QuantumSupernet, all while incurring lower quantum computational costs. Notably, this task involves a much larger search space than any previously examined task, and the TD-QAS framework demonstrates significant improvements in QAS performance compared to earlier Numerical simulations. These results further emphasize the importance of the TD-QAS framework in effectively reducing the search space and improving quantum circuit optimization.
@@ 4.3.2 TD-DQAS framework
To validate the generality and effectiveness of the TD-QAS framework across different QAS algorithms, this section implements the TD-QAS framework using the DQAS algorithm and conducts a systematic performance analysis. The core algorithm of DQAS has been detailed in Section 2.1.2. Similar to Section 4.3.1, we first describe the implementation details of the TD-QAS framework using DQAS, followed by an introduction to the hyperparameter settings. Finally, we execute the QAS algorithms and analyze their results.
Overview of Framework
In this section, we outline the process of implementing the TD-QAS framework with DQAS. The goal of DQAS is to identify a high-performance quantum circuit that contains $N_{\text{gate}}$ quantum gates, where arbitrary single-qubit or two-qubit gates can be placed at any position. In the TD phase, the native gate set is defined as $A_{\text{t}}$ is $\{R_{x},XX\}$. DQAS uses a probabilistic model to select a batch of $N_{\text{batch}}$ quantum circuits, each of which is composed of $N_{\text{gate}}$ gates chosen from the set $A_{\text{t}}$ and their corresponding qubit positions. The performance of the selected circuits is evaluated using shared parameters, which are then updated using feedback from the quantum circuit evaluation. This process continues until $N_{\text{train}}$ iterations or until convergence. Finally, the probability model selects the optimal topology instance with the highest probability. DQAS uses mini-batch gradient descent to optimize the model. Generally, models with greater complexity require a larger batch size. In the GT phase, the shared parameters from the TD phase are inherited, and the probability model is trained from scratch. The model fine-tunes the quantum gates from the native gate set based on the searched topology. Since gate placement positions are not considered in this phase, the trainable parameters in the probability model are minimal, allowing for a smaller $N_{\text{train}}$.
@@
Table 4: Numerical Simulation Results for TD-DQAS.
@@
This section validates the capacity of the TD-DQAS framework across three tasks: the VQE task, the MaxCut problem, and the state classification task. Since the numerical simulation setup and conclusions largely overlap with those discussed in earlier sections, we focus on presenting a comparative analysis of the results rather than repeating the full details.
For each task, we set the following hyperparameters: $A=A_{\text{g}}=\{R_{x},R_{y},R_{z},XX,YY,ZZ\}$ $A_{\text{t}}=\{R_{x},XX\}$. $N_{\text{train}}=500$, the $N_{\text{batch}}$ of DQAS, TD and GT phase is 32, 8, 8, respectively. For the VQE task, we selected the 6-qubit TFIM, using the ground state energy as the performance metric. We set $N_{\text{qubit}}=6$ and $N_{\text{gate}}=36$. Therefore, the search space of DQAS, TD and GT phase is $36^{36}$, $12^{36}$ and $3^{36}$. The search space reduction ratio is nearly $1.5\times 10^{17}$. In the MaxCut problem, we used the ER model to randomly generate 100 distinct graphs, each containing 10 nodes, with an edge creation probability set to 0.5. The quantum circuit’s output apart from the maximum cut, serves as the performance metric. We set $N_{\text{qubit}}=10$ , $N_{\text{gate}}=20$. Therefore, the search space of DQAS, TD and GTphase is $60^{20}$, $20^{20}$ and $3^{20}$. The search space reduction ratio is nearly $3.5\times 10^{10}$. For the quantum state classification task, classification accuracy serving as the performance metric. We set $N_{\text{qubit}}=8$ , $N_{\text{gate}}=30$. Therefore, the search space of DQAS, TD, and GT phase is $48^{30}$, $16^{30}$ and $3^{30}$. The search space reduction ratio is nearly $2\times 10^{15}$. The detailed numerical results for each task are presented in Table 4, which shows that the TD-DQAS framework outperforms the DQAS algorithm across all tasks. In both noiseless and noisy scenarios, TD-QAS significantly enhances the performance of QAS algorithms, enabling them to search for higher-performance quantum circuits with lower quantum resource costs.

### v2

@@ 4.3.1 TD-QuantumSupernet framework
In this subsection, we detail the implementation of the TD-QAS framework using QuantumSupernet, followed by a series of numerical simulations. First, we explain how to implement the TD and GT phases using QuantumSupernet. The core algorithm of QuantumSupernet is described in Section 2.1.1. Subsequently, we define the hyperparameters for various numerical simulation tasks, highlighting their impact on the comparison variables. Finally, we set the hyperparameters, execute the algorithm, and analyze the results of the numerical simulations. For the simulations, we select the same VQE and classification tasks used in QuantumSupernet.
Implementation of TD-QuantumSupernet
@@
Here, we describe the implementation of the TD-QAS framework within the QuantumSupernet algorithm. The TD phase begins by defining the topology search space based on the supernet structure and the native gate set. The number of qubits ($N_{\text{qubit}}$) and the number of layers ($N_{\text{layer}}$) are specified by the task. Each layer supports up to $N_{\text{qubit}}$ quantum gates. The native gate set is defined as $\{R_{x},I,CNOT,CI\}$, where $CI$ refers to the Controlled-Identity Gate. If a sampled topology includes an $I$ or $CI$ gate at a specific location, no quantum gate will be placed there, and the GT phase will bypass fine-tuning for that position. Subsequently, we utilize random search strategy to explore potential high-performance topologies. Each quantum circuit is evaluated using shared parameters, with feedback guiding the updates to these parameters. Specifically, $N_{\text{experts}}$ shared parameters are used to assess the performance of each topology, allowing for more precise evaluations. During the initial $T_{\text{warm}}$ iterations, a randomly selected expert from the $N_{\text{experts}}$ pool evaluates the topology instantiation and updates the shared parameters accordingly. Following the warm-up phase, all $N_{\text{experts}}$ contribute to the evaluation, but only the parameters with the best performance are updated. When the shared parameters have converged, we evaluate $N_{\text{search}}$ topology instantiations, selecting the optimal topology. In the gate-type (GT) phase, the search space is determined by the selected topology and the gate set $A_{\text{g}}$. To expedite training, the shared parameters from the TD phase are directly inherited and adjusted to accommodate the new gate configuration. Specifically, this inheritance mechanism involves replicating the shared parameters from the TD phase based on the number of single-qubit and two-qubit gates available in the native gate pool. To enhance the accuracy of the evaluation, we performed additional training on the inherited parameters for a few extra iterations, denoted as $T_{\text{extra}}$, which further refines the assessment of specific circuit performances.
@@
Based on the defined hyperparameters, we compute key comparison variables, including the size of the search space and the total quantum computational cost(QCC). As the supernet consistently applies both single-qubit and two-qubit gates across all qubits at each layer, the original search space is calculated as $(|A[\text{single}]|^{N_{\text{qubit}}}\cdot|A[\text{double}]|^{N_{\text{qubit}}})^{N_{\text{layer}}}$, where $|A[\text{single}]|$ denotes the number of single-qubit gates in the native gate set. The topology search space is defined as $(2^{N_{\text{qubit}}})^{N_{\text{layer}}}$, where $2^{N_{\text{qubit}}}$ accounts for all possible combinations of single- and two-qubit placeholders per layer. The gate-type search space, conditioned on a specific topology, is expressed as $|A[\text{single}]|^{N_{\text{layer}}-\text{x}}\cdot|A[\text{double}]|^{\text{x}}$, where $x$ denotes the number of double-qubit placeholders in the selected topology. The quantum computational cost (QCC) for QuantumSupernet and the TD phase is given by $\sum_{\text{1}}^{T_{\text{warm}}}T_{\text{i}}+\sum_{T_{\text{warm}+1}}^{T}T_{\text{i}}*N_{\text{experts}}+\sum_{\text{1}}^{N_{\text{search}}}T_{\text{i}}*N_{\text{experts}}$ The QCC for the GT phase is denoted by $\sum_{\text{1}}^{N_{\text{search}}}T_{\text{i}}\cdot(1+T_{\text{extra}})$. Here, $T_{i}$ represents the quantum computational cost of a single circuit evaluation, as defined in Equation 7.
@@
In this section, we evaluate the performance of the TD-QAS framework through numerical simulations. First, we replicate the ground state energy estimation task for a 4-qubit hydrogen molecule, which is consistent with the simulation in Ref. [23]. While this small-scale task provides valuable insight, it does not fully demonstrate the potential of the TD-QAS framework. Therefore, we extend the validation to larger tasks, including the 5-qubit Heisenberg model.
For the simulation task estimating the ground state energy of hydrogen (which has a ground state energy of -1.13618 Hartree), the Hamiltonian construction is detailed in QuantumSupernet [23]. Before performing the numerical simulations, we define the hyperparameters and compare the search space sizes of the QuantumSupernet and TD-QAS methods. Specifically, we set the number of qubits ($N_{\text{qubit}}$) to 4 and the number of layers ($N_{\text{layer}}$) to 3. For the QuantumSupernet, the native gate set $A$ is defined as $\{I,R_{y},R_{z},CNOT,CI\}$. In the TD-QAS framework, the gate sets are defined as $A_{\text{t}}=\{R_{y},I,CNOT,CI\}$ for the topology-driven phase and $A_{\text{g}}=\{R_{y},R_{z},CNOT\}$ for the gate tuning phase. Under these settings, the search space for the QuantumSupernet is calculated as $(3^{4}\cdot 2^{4})^{3}$. The search space for the TD phase is calculated as $(2^{4}\cdot 2^{4})^{3}$, while the GT phase operates within a search space ranging from 1 to $2^{12}$. This decoupling yields a search space reduction ratio of $2.5\times 10^{1}$ compared to the original QuantumSupernet. Owing to the reduced search space, we employ smaller values for the number of experts ($N_{\text{experts}}$) and the total number of iterations ($T$) to enhance efficiency. Specifically, for the QuantumSupernet, we set $N_{\text{experts}}=5$, $T=500$, $T_{\text{warm}}=200$, and $N_{\text{search}}=500$. In the TD phase, these parameters are adjusted to $N_{\text{experts}}=1$, $T=200$, $T_{\text{warm}}=100$, and $N_{\text{search}}=500$. In the GT phase, we set $T_{\text{extra}}=1$. Consider the small problem size and the limited size of the native gate set, an exhaustive search is feasible in the GT phase, thus we omit $N_{\text{search}}$ for this stage.
@@
To further demonstrate the capabilities of the TD-QAS framework for larger-scale problems, we applied it to the ground state energy estimation of the 5-qubit Heisenberg model. The Hamiltonian for the Heisenberg model is defined as $\sum_{\text{i=1}}^{\text{n}}\left(X_{\text{i}}X_{\text{i+1}}+Y_{\text{i}}Y_{\text{i+1}}+Z_{\text{i}}Z_{\text{i+1}}\right)+\sum_{\text{i=1}}^{\text{n}}Z_{\text{i}}$ , and the ground state energy is -8.47213. The hyperparameters for this simulation are defined as $N_{\text{qubit}}$=5, $N_{\text{layer}}$= 6, $A$={$I$, $R_{x}$, $R_{y}$, $R_{z}$, $CNOT$, $CI$}, $A_{\text{t}}$={$R_{y}$, $I$, $CNOT$, $CI$}, and $A_{\text{g}}$={$R_{x}$, $R_{y}$, $R_{z}$, $CNOT$}. With these settings, the search space for the QuantumSupernet is calculated as $(4^{5}\cdot 2^{5})^{6}=2^{80}$, while the search space for the TD and GT phases is $2^{60}$ and $1\sim 3^{30}$, respectively. This results in a search space reduction ratio of approximately $1\times 10^{7}$. To evaluate the performance of the TD-QAS framework and compare it to the QuantumSupernet, we executed simulations with different values of $N_{\text{experts}}$. Specifically, we set $N_{\text{experts}}=1$ and $N_{\text{experts}}=5$. The hyperparameters for the QuantumSupernet are $T=500$, $T_{\text{warm}}=200$, and $N_{\text{search}}=500$. For the TD-QAS framework, the parameters are adjusted as $T=200$, $T_{\text{warm}}=100$, and $N_{\text{search}}=300$. In the GT phase, we set $T_{\text{extra}}=1$ and $N_{\text{search}}=100$.
@@
The results of the numerical simulations are presented in Table 2. With the same $N_{\text{experts}}$, the TD and GT methods outperform the QuantumSupernet. Therefore, the TD-QAS framework enhances the performance of the QuantumSupernet. When comparing simulations with different numbers of experts ($N_{\text{experts}}=1$ and $N_{\text{experts}}=5$), we observe that the QuantumSupernet shows a higher dependency on $N_{\text{experts}}$. This is owing to various parameterized quantum gates in the circuits searched by QuantumSupernet, which results in greater sensitivity to the number of experts. However, the circuits found by the TD and GT methods have fewer parameters, resulting in more stable performance across varying values of $N_{\text{experts}}$. This highlights the TD-QAS framework’s ability to reduce the training burden, which in turn optimizes the QCC. Additionally, the circuits discovered in both the TD and GT phases exist within the operation search space, but were not identified by the QuantumSupernet. Therefore, the TD-QAS framework may help mitigate the risk of the QAS algorithm becoming trapped in local optima, enhancing the discovery of more optimal quantum circuits.
@@
To assess the performance improvements of the TD-QAS framework on more complex tasks, we conducted a numerical evaluation using an 8-qubit quantum state classification task. Specifically, we employed a QNN to address a binary classification problem, where the input comprised quantum states with varying levels of entanglement, characterized by Concurrence Entropy (CE = 0.15 and CE = 0.45). The quantum dataset used in this experiment is based on the dataset from Ref. [46].
The hyperparameter settings for the numerical simulation are detailed as follows, $N_{\text{qubit}}$=8, $N_{\text{layer}}$= 6, $A$={$I$, $R_{x}$, $R_{y}$, $R_{z}$, $CNOT$, $CI$}, $A_{\text{t}}$={$R_{y}$, $I$, $CNOT$, $CI$}, and $A_{\text{g}}$={$R_{x}$, $R_{y}$, $R_{z}$, $CNOT$}. With these parameters, we calculate that the search space for the QuantumSupernet is $2^{144}$, while the search space for the TD and GT phases is $2^{96}$ and $1\sim 3^{30}$, respectively. This results in a search space reduction ratio of approximately $2.6\times 10^{6}$. The simulation hyperparameters for the QuantumSupernet are set as $T=500$, $T_{\text{warm}}=200$, and $N_{\text{search}}=500$. For the TD-QAS framework, we set $T=200$, $T_{\text{warm}}=100$, and $N_{\text{search}}=300$ in the TD phase, while in the GT phase, $T_{\text{extra}}=1$ and $N_{\text{search}}=100$.
Table 3 summarizes the simulation results for the state classification task. The TD-QAS framework successfully identifies higher-performance circuits compared to the QuantumSupernet, all while incurring lower quantum computational costs. Notably, this task presents a larger search space than any previously examined, and the TD-QAS framework significantly enhance QAS performance compared to earlier Numerical simulations. These results further emphasize the importance of the TD-QAS framework in effectively reducing the search space and enhancing search optimization.
@@ 4.3.2 TD-DQAS framework
To validate TD-QAS framework’s generality and effectiveness across different QAS algorithms, we implement it using the DQAS algorithm and systematically analyze its performance. The core algorithm of DQAS has been detailed in Section 2.1.2. Similar to Section 4.3.1, we first describe the implementation details of the TD-QAS framework using DQAS, followed by an introduction to the hyperparameter settings. Finally, we execute the QAS algorithms and analyze their results.
Implementation of the TD-DQAS
Here, we outline the process of implementing the TD-QAS framework with DQAS. DQAS aims to identify a high-performance quantum circuit that contains $N_{\text{gate}}$ quantum gates, where arbitrary quantum gates can be placed at any position. In the TD phase, the native gate set is defined as $A_{\text{t}}$ is $\{R_{x},XX\}$. DQAS employs a search strategy base on probabilistic model to select a batch of $N_{\text{batch}}$ quantum circuits, each of which is composed of $N_{\text{gate}}$ gates chosen from the set $A_{\text{t}}$ and their corresponding qubit positions. The performance of the selected circuits is evaluated using shared parameters, which are updated by feedback from the quantum circuit evaluation. This process continues until $N_{\text{train}}$ iterations or when it converge. Subsequently, the probability model selects the optimal topology with the highest probability. DQAS employs mini-batch gradient descent to optimize the model. Generally, models with greater complexity require a larger batch size. In the GT phase, the shared parameters from the TD phase are inherited, and the probability model is trained from scratch. The model fine-tunes the quantum gates from the native gate set based on the searched topology. As the probability model only should consider gate types in the GT phase, the number of trainable parameters is small, allowing for a smaller $N_{\text{train}}$.
@@
Table 4: Numerical simulation results for TD-DQAS on three tasks. The ”Performance” column shows the output metric for each task (energy, approximation ratio, or accuracy). QCC denotes quantum computational cost.
@@
This section validates the capacity of the TD-DQAS framework across three tasks: the VQE task, the MaxCut problem, and the state classification task. As numerical simulation setup and conclusions largely overlap with those discussed in earlier sections, we focus on presenting a comparative analysis of the results rather than repeating the full details.
For each task, we set the following hyperparameters, $A=A_{\text{g}}=\{R_{x},R_{y},R_{z},XX,YY,ZZ\}$ $A_{\text{t}}=\{R_{x},XX\}$. $N_{\text{train}}=500$, the $N_{\text{batch}}$ of DQAS, TD and GT phase is 32, 8, 8, respectively. For the VQE task, we selected the 6-qubit TFIM, using the ground state energy as the performance metric. We set $N_{\text{qubit}}=6$ and $N_{\text{gate}}=36$. Therefore, the search space of DQAS, TD and GT phase is $36^{36}$, $12^{36}$ and $3^{36}$. The search space reduction ratio is nearly $1.5\times 10^{17}$. In the MaxCut problem, we employed the ER model to randomly generate 100 distinct graphs, each containing 10 nodes, with an edge creation probability set to 0.5. The approximation ratio serves as the performance metric. We set $N_{\text{qubit}}=10$ , $N_{\text{gate}}=20$. Therefore, the search space of DQAS, TD and GTphase is $60^{20}$, $20^{20}$ and $3^{20}$. The search space reduction ratio is nearly $3.5\times 10^{10}$. For the quantum state classification task, classification accuracy serving as the performance metric. We set $N_{\text{qubit}}=8$ , $N_{\text{gate}}=30$. Therefore, the search space of DQAS, TD, and GT phase is $48^{30}$, $16^{30}$ and $3^{30}$. The search space reduction ratio is nearly $2\times 10^{15}$.
Table 4 presents detailed numerical results for each task, demonstrating that the TD-DQAS framework outperforms the DQAS algorithm across all tasks. Specifically, in the VQE, MaxCut, and quantum state classification tasks, the TD-QAS framework achieves higher performance with significantly reduced quantum computational costs. Therefore, decoupling the search process into two phases—topology search and gate-type fine-tuning—improves the efficiency of QAS algorithms, enabling them to obtain higher-performing quantum circuits while minimizing resource requirements.